# When Respondents Don’t Care Anymore: Identifying the Onset of Careless Responding

### v2

@@ Introduction
For each participant, we can obtain multiple individual series of length equal to the total number of items in the survey. Each individual series should thereby measure one of the dimensions indicative of carelessness. Although we focus on using reconstruction errors from (1) to measure inconsistency as well as the LSP sequence from (2) to measure invariability, it should be noted that in principle CODERS can also be used with any single measurement or any other combination of different measurements of carelessness.
@@
As motivated in the previous two sections, we expect a changepoint in any of the measurements of carelessness from the onset of careless responding (e.g., Figure 2(a)), and no changepoint in the absence of carelessness. We therefore propose to apply the nonparametric cumulative sum self-normalization test of Shao & Zhang, (2010), which was developed in the context of time series analysis. It searches for the location of a possible changepoint in a multivariate series and has attractive theoretical guarantees (see Shao & Zhang,, 2010; Zhao et al.,, 2022, and our detailed description in Appendix B.3).
@@
For each point in a multivariate series (we use measurements of carelessness associated with each questionnaire item), this changepoint detection test calculates a statistic that compares a quantity of interest (we take the mean) in the periods before and after the current point (item). If the maximum of this statistic exceeds a critical value, the point (item) where this maximum occurs is flagged as a changepoint. Otherwise no changepoint is flagged. The critical value thereby depends only on the significance level $\alpha$ and the dimension of the multivariate series. Importantly, it controls the size of the test regardless of the distribution of the individual series. This gives CODERS great flexibility, as it allows us to use any combination of measurements of carelessness without making possibly unrealistic assumptions on their distributions or obtaining cutoff values in an ad-hoc manner. In line with literature that suggests that there should be overwhelming evidence when labeling respondents as careless (e.g., Huang et al.,, 2012), we recommend using a conservative significance level no greater than 1% for identifying changepoints.
@@
As an example, consider again the two-dimensional series in Figure 2(a), which measures carelessness of a certain survey participant in terms of autoencoder reconstruction errors (RE) and the LongstringPattern (LSP) sequence. Figure 2(b) shows the corresponding series of statistics, where the maximum value at the 148th item clearly exceeds the critical value. That is, CODERS flags this respondent as being careless from the 148th item onwards. Indeed, from roughly this item onwards, the reconstruction errors are in general much larger than before. It follows that this respondent likely is a partially inconsistent careless respondent.
@@
Ideally, new methodology is empirically validated by conducting an experiment that emulates the phenomenon of interest in lab conditions. For instance, Schroeders et al., (2022) conduct an experiment in which they instruct one part of the participants to respond truthfully and another part to respond carelessly in order to investigate whether various screeners accurately identify the subset of careless respondents. For any of the compared methods, at most 19% of the respondents flagged as careless were in fact instructed to be careless. The authors therefore express doubts whether the study participants complied with the experimental instructions. In addition to the risk of non-compliance, Ulitzsch et al., 2022a ; Ulitzsch et al., 2022b stress that another major limitation of this experimental approach is the doubtful assumption that participants instructed to respond carelessly behave in a comparable manner to careless respondents outside of such lab experiments. Indeed, it seems plausible that careless responding stems from an intrinsic state-of-mind rather than being a conscious action that can be instructed (cf. Alfons & Welz,, 2024, for additional discussion). Hence, rather than conducting such an experiment, we take data from a seminal study on careless responding and further examine the responses for partial carelessness.
@@
We apply CODERS to an empirical data set collected by Johnson, (2005). The data concern the five factor model of personality (Goldberg,, 1992), which assumes that variation in a measurement of personality can be explained by the so-called Big Five personality traits (factors): openness, conscientiousness, extraversion, agreeableness, and neuroticism. Each of the Big Five factors can be further decomposed into six subdomains (facets): for instance, depression and anxiety are facets of the neuroticism trait. Consquently, common Big Five instruments, such as the 240-item NEO-PI-R (Costa & McCrae,, 1992), typically measure a total of $5\times 6=30$ facets. The data set of Johnson, (2005) consists of internet-collected responses to a 300-item representation of the NEO-PI-R inventory (Goldberg,, 1999), in which each of the 30 facets is measured by 10 items on a five-point Likert scale.
@@
In general, it seems likely that survey fatigue is present in a sample of responses to an online questionnaire with a lengthy battery of 300 items. For instance, Bowling et al., (2021) estimate that if one wishes that at least 90% of participants do not respond carelessly to more than 20% of all items, one should not include more than 176 items in an online questionnaire. For the study of Johnson, (2005), it is important to keep in mind that the author applied two methods for detecting inconsistent careless respondents to the data from which straightliners are already removed. The two methods are designed to detect participants who respond carelessly to all or at least the vast majority of survey items. It is unclear how well they perform in identifying partially careless respondents, who may respond carelessly to only a relatively small number of items. In addition, Johnson, (2005) did not apply any method designed to identify invariable careless responding other than straightlining. By applying CODERS, which is designed to identify partially careless responding for various types of carelessness, new insights into the incidence and nature of (partially) careless responding may emerge.
@@ Results
Comparison with Benchmark Methods
For obtaining the autoencoder reconstruction errors in CODERS, we set the number of nodes in the bottleneck layer to 30, i.e., equal the number of facets in the Big Five model. Table 1 summarizes the results of CODERS. Depending on the significance level $\alpha$, CODERS flags between 164 (0.7%) participants for $\alpha=0.001$ and 850 (3.8%) participants for $\alpha=0.01$ as partially careless. On average, the location of the carelessness onset item for flagged participants is estimated around the 150th item, meaning that carelessness onsets on average after half of the 300 total items. These findings are broadly in line with the results of Bowling et al., (2021, Figure 1), whose model suggests a low probability that a representative participant starts responding carelessly at the 150th item. The earliest onset occurs at item 12, whereas the latest onset is at item 294. A more detailed visual overview of the detected onset items for the flagged participants can be found in Figure C.1 in Appendix C.2.
For comparison, we consider personal reliability and psychometric antonym scores. Along the given responses, Johnson, (2005) provides per-individual estimates for either score, which we use in our analysis. Importantly, Johnson, reverses psychometric antonym scores so that low scores are associated with carelessness in both methods. Following Johnson,, we flag respondents as careless whose personal reliability and psychometric antonym scores are lower than $0.3$ and $-0.03$, respectively. The results are shown in Table 1. Personal reliability flags 40 participants (0.2%) as careless, while psychometric antonym flags 186 respondents (0.8%). CODERS, personal reliability, and psychometric antonyms generally show little agreement regarding the detected careless respondents. Personal reliability and psychometric antonym scores correlate by a moderate value of 0.488, yielding only 13 participant that are flagged by both methods. Depending on the significance level $\alpha$, CODERS and personal reliability have between 1 and 4 flagged participants in common, while the number of jointly flagged respondents for CODERS and psychometric antonym ranges between 5 and 10. All three methods share a single flagged participant.
Table 1: Comparison of CODERS with benchmark methods for the data of Johnson, (2005).
@@
Reliability 40 (0.2%) 1 3 4
Antonym 186 (0.8%) 5 7 10
Both 13 1 1 1
@@
It is not surprising that there is little to no consensus of CODERS with the two benchmark methods, as the former is designed for a different task: finding respondents that have been careless in parts of the questionnaire (including identifying the onset item) rather than throughout the entire questionnaire. This disagreement indicates that partial carelessness may be present in the data, which the benchmark methods may be unable to identify. We therefore take a closer look at a representative selection of respondents and analyze their per-item reconstruction errors and LSP sequences, which are shown in Figure 3. Table 2 contains additional summary statistics for these measurements before and after the estimated careless onset item (if such an item is found by CODERS). Two of the selected participants are flagged as careless by CODERS but not by either of the two benchmark methods, one participant is flagged by both CODERS and psychometric antonym, and one participant is flagged by both benchmark methods but not CODERS.
Figure 3: Scores and careless onset item detected by CODERS in selected participants from the data of Johnson, (2005).
(a) Flagged by CODERS only, presumably inconsistent. Onset is estimated at item 148.
(b) Flagged by CODERS only, presumably invariable. Onset is estimated at item 58.
(c) Flagged by antonym and CODERS, presumably inconsistent. CODERS estimates onset at item 119.
(d) Flagged by antonym and personal reliability, but not CODERS (i.e., there is no estimated onset item).
Note. In each subplot, autoencoder reconstruction errors (RE) and LongstringPattern (LSP) sequences are shown in separate panels. The solid vertical lines indicate the identified careless onset items. Summary statistics for the corresponding participants are provided in Table 2.
Table 2: Average scores before and after the careless onset item detected by CODERS in selected participants from the data of Johnson, (2005).
@@
LSP 4.197 (1.396) 3.497 (1.433)
(a) Flagged by CODERS only, presumably inconsistent. Onset is estimated at item 148.
@@
LSP 2.421 (0.706) 19.770 (10.941)
(b) Flagged by CODERS only, presumably invariable. Onset is estimated at item 58.
@@
RE 0.019 (0.022) 0.039 (0.059)
LSP 3.619 (1.371) 2.687 (0.972)
(c) Flagged by antonym and CODERS, presumably inconsistent. CODERS estimates onset at item 119.

### v3

@@ Introduction
For each participant, we can obtain multiple individual series of length equal to the total number of items in the survey. Each individual series should thereby measure one of the dimensions indicative of carelessness. Although we focus on using reconstruction errors from (1) to measure inconsistency as well as the LSP sequence from (2) to measure invariability, in principle CODERS can also be used with any single measurement or any other combination of different measurements of carelessness on the participant-item level.
@@
As motivated in the previous two sections, we expect a changepoint in any of the measurements of carelessness from the onset of careless responding (cf. Figures 2 and 3), and no changepoint in the absence of carelessness. We therefore propose to apply the nonparametric cumulative sum self-normalization test of Shao & Zhang, (2010), which was developed in the context of time series analysis. It searches for the location of a possible changepoint in a multivariate series and has attractive theoretical guarantees (see Shao & Zhang,, 2010; Zhao et al.,, 2022, and our detailed description in Appendix B.4).
@@
For each point in a multivariate series (we use measurements of carelessness associated with each questionnaire item), this changepoint detection test calculates a statistic that compares a quantity of interest (we take the mean) in the periods before and after the current point (item). If the maximum of this statistic exceeds a critical value, the point (item) where this maximum occurs is flagged as a changepoint. Otherwise no changepoint is flagged. The critical value thereby depends only on a pre-specified significance level $\alpha$ and the dimension of the multivariate series. Importantly, it controls the size of the test regardless of the distribution of the individual series. This gives CODERS great flexibility, as it allows us to use any combination of measurements of carelessness without making possibly unrealistic assumptions on their distributions or obtaining cutoff values in an ad-hoc manner. In line with literature that suggests that there should be overwhelming evidence when labeling respondents as careless (e.g., Huang et al.,, 2012), we recommend using a conservative significance level no greater than 1% for identifying changepoints.
@@
As an illustrative example, consider Figure 4(a). It illustrates two dimensions computed from a given participant’s responses, namely autoencoder reconstruction errors and the LSP sequence. The former series has been shown before in Figure 2. The bottom panel of Figure 4(a) shows the test statistics computed from this two-dimensional series, where the maximum value at the 172nd item clearly exceeds the critical value. That is, CODERS identifies inconsistent carelessness from the 172nd item onward, as the reconstruction errors are in general much larger than before. Indeed, this simulated participant started choosing responses at random at the identified item. Figure 4(b) is similar, but computed from a different respondent, whose LSP sequence has been shown before in Figure 3. CODERS flags a changepoint at the 261st item. From this item onward, the LSP values are clearly elevated, indicating invariable carelessness. In fact, this simulated participant started repeating the pattern $3-4$ just one item before the estimated onset.
Figure 4: Illustrative example output from CODERS using responses of two simulated partially careless respondents to a 300-item questionnaire.
(a) Inconsistent partial carelessness.
(b) Invariable partial carelessness.
Note. The left subfigure shows the output from CODERS for a simulated partially inconsistent respondent who starts responding randomly from the 172nd item onward, while the right subfigure does so for a simulated partially invariable respondent who starts repeating the pattern $3-4$ from the 260th item onward. Each subfigure visualizes the autoencoder reconstruction errors (RE), LongStringPattern (LSP) sequence, and test statistics associated with the corresponding participant’s responses. The blue solid vertical lines indicate the carelessness onset items identified by CODERS at significance level 0.1%, which correspond to the true onset item 172 for the partially inconsistent respondent, and to item 261 (one item after true onset) for the partially invariable respondent. The dashed horizontal line is the critical value of the test statistic at the 0.1% significance level (192.5; see Table B.2 in Appendix B.4). The reconstruction errors in the left subfigure are those shown in Figure 2, and the LSP sequence in the right subfigure is that shown in Figure 3.
Simulation Experiments
After introducing CODERS, we now investigate its behavior via simulation experiments modeled after data collected by Johnson, (2005). We use the data from this seminal study on careless responding in our empirical application in a later section, where the data are described in more detail.
Data Generation
The survey instrument used by Johnson, (2005) is a 300-item representation of the NEO-PI-R (Goldberg,, 1999) measurement of the Big Five personality traits (openness, conscientiousness, extraversion, agreeableness, and neuroticism; Goldberg,, 1992) by means of five-point Likert scales. Each of the five traits comprises six subdomains (facets) so that a total of $5\times 6=30$ facets are measured. We base our simulation design on the between-item correlation structure and per-item response distributions in the data collected by Johnson, (2005), calculated after removing any presumably careless respondents detected in our empirical application. Figure 5 visualizes the resulting correlation matrix and marginal response distributions. Of the 300 items, 152 are positively scored while 148 are negatively scored. In the correlation matrix, one can identify the five traits and, within each block of within-trait correlations, also the six corresponding facets. Using these correlations and marginal response distributions, we generate rating-scale responses of $n=500$ participants to the $p=300$ items,44endnote: 4We use package simstudy (Goldfeld & Wujciak-Jens,, 2020) to generate the data, for which the target values of correlations only hold approximately on the population level. with the items in the same randomized order as in Johnson, (2005).
Figure 5: Correlation matrix and marginal response distributions used in simulations.
(a) Correlation matrix of the $p=300$ items. A more saturated blue (red) in a cell corresponds to a stronger positive (negative) correlation between the corresponding items.
(b) Marginal response distributions of the $p=300$ items ($x$-axis), each of which has five response options ($y$-axis). A darker shade of gray indicates a higher probability of the corresponding response option.
Note. The correlation matrix and the marginal response distributions are calculated from the data of Johnson, (2005) after removing presumably careless respondents. Of the $p=300$ items, 152 are positively scored and 148 are negatively scored. The items in both plots are presented in an order such that items measuring the same facet are adjacent.
Subsequently, a fixed percentage $\gamma\in\{0,0.02,0.04,0.1,0.2,0.3,0.4,0.5,0.6,0.8,1\}$ of respondents is selected to be partially careless. That is, $\gamma$ denotes the prevalence of careless responding. For each of the selected partially careless respondents, we sample a carelessness onset item from which onward the remaining responses are careless.
We consider three regimes for the carelessness onset: (i) the baseline setting where the onset item is randomly sampled from $\{30,31,\dots,269,270\}$ so that carelessness starts between 10% and 90% of all items, (ii) early onset randomly sampled from items $\{30,31,\dots,149,150\}$ (between 10% and 50% of all items), and (iii) late onset randomly sampled from items $\{150,151,\dots,269,270\}$ (between 50% and 90% of all items).
[…]
We repeat the above described data generating process 100 times. We compare three variations of CODERS based on (i) both reconstruction errors (RE) and the LSP sequence, (ii) RE only, and (iii) the LSP sequence only. For the autoencoder, we set the number of nodes in the bottleneck layer equal to the number of measured constructs (i.e., 30 facets). In all three variations of CODERS, we look for changepoints with three different significance levels $\alpha$, namely 0.1%, 0.5%, and 1% (from most to least conservative).
Performance Evaluation
To quantify how accurately partially careless respondents are detected, we compute the false positive rate (FPR) (proportion of true attentive respondents incorrectly flagged as careless) and the false negative rate (FNR) (proportion of true careless respondents incorrectly not flagged as careless). For CODERS, we consider a respondent to be flagged as careless if a changepoint is detected. In addition, to quantify how accurately the onset of careless responding is detected, we compute the mean absolute error (MAE) (difference between the true onset item and the detected onset item in absolute value, averaged over the set of true careless respondents for whom a changepoint is detected). We report averages of these performance measures across the 100 simulation runs. For all three measures, a lower value (close to 0) indicates better performance.
Results
We focus on discussing the results for CODERS using both measurements of carelessness (RE and LSP), except when we explicitly state otherwise. Varying the prevalence of careless responding, Figure 6 visualizes the FPR of CODERS in fully attentive respondents for the different carelessness onset regimes (in separate columns). Overall, CODERS rarely flags changepoints in fully attentive respondents. Even with our most liberal choice of significance level $\alpha$ (1%), overrejection increases only slowly to about 3% for 60% prevalence, before it reaches between 3.5% and 5.7% at 80% prevalence (depending on the onset regime). Other than for extremely high prevalence levels, the FPR stays reasonably close to the respective nominal level, indicating that CODERS controls Type I errors fairly well.
Figure 6: False positive rate of CODERS for attentive respondents.
Note. The $x$-axis displays various prevalence levels of partially careless respondents. Results for different onset regimes are shown in separate columns, averaged across 100 repetitions.
For partially careless respondents, Figure 7 (top row) contains a similar plot of the FNR. Most importantly, the FNR of CODERS with both dimensions remains remarkably constant throughout prevalence levels, at about 18–26% in the baseline onset regime depending on the significance level. While the FNR increases to about 36% with the most conservative significance level (0.1%) in the early onset regime, it decreases to about 8% with the most liberal significance level (1%) in the late onset regime. Overall, the relatively low FNR implies that CODERS successfully identifies a large majority of the true careless respondents. Combined with its low FPR, we find that CODERS discriminates well between attentive and partially careless respondents.
In addition, Figure 7 (bottom row) displays the MAE of the estimated onset in partially careless respondents. The prevalence level and onset regime have little effect on the accuracy of the estimates, with an MAE hovering around 2 to 3 depending on the significance level. That is, on average the estimated onset is only two to three items away from the true onset, indicating that CODERS achieves excellent accuracy in estimating carelessness onset.
Figure 7: Performance measures of CODERS for partially careless respondents.
Note. The $x$-axis displays various prevalence levels. The top row shows results for the false negative rate, while the bottom row visualizes the accuracy of onset estimation via the mean absolute error. Results for different onset regimes are shown in separate columns, averaged across 100 repetitions.
On the other hand, Figure 7 further reveals that the FNR of CODERS using only RE or LSP is considerably higher in all settings. Furthermore, the MAE of the estimated onset is much higher when using only RE. It is important to note that such a drop in performance is expected when using only one measurement of carelessness due to the various types of careless responding being simulated. RE is designed to capture inconsistent carelessness and may therefore miss invariable carelessness, whereas it is the other way around for LSP. To investigate this further, Figures 8 and 9 show the FNR and the MAE of onset estimation averaged over subgroups of participants corresponding to each type of careless responding (in separate rows). For the two inconsistent types of carelessness (top two rows), using both dimensions yields overall good results similar to to those using only RE, whereas performance is clearly worse when using only LSP. For the two invariable types of carelessness (bottom two rows), the results when using both dimensions are near-perfect (with FNR and MAE close to 0) and indistinguishable from using only LSP, while performance deteriorates when using only RE. Hence, these findings highlight the benefit of combining measurements of different dimensions of carelessness in CODERS.
Figure 8: False negative rate of CODERS for different types of partially careless respondents.
Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding (rows) in different onset regimes (columns) are averaged across 100 repetitions.
Figure 9: Mean absolute error of carelessness onset estimation with CODERS for different types of partially careless respondents.
Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding (rows) in different onset regimes (columns) are averaged across 100 repetitions.
We now zoom in on the results of CODERS for the two types of inconsistent carelessness in Figures 8 and 9, as there is notable heterogeneity. Specifically, CODERS performs clearly better for extreme responding than for random responding, with the FNR and the MAE of onset estimation in general being lower. Considering that the items exhibit a variety of different probability distributions for the attentive responses, with a low probability mass in at least one of the extreme answer categories (see Figure 5), distinguishing random responding is simply a much more challenging problem than distinguishing extreme responding.
For random responding (see the first row in Figure 8), CODERS only achieves a decent FNR in the late onset regime with significance level 1% or 0.5% (about 32% and 40%, respectively). In the baseline onset regime, its FNR is higher than 50%. In the early onset regime, even our most liberal significance level of 1% only yields an FNR of about 75%. Nevertheless, this finding should not be mistaken as an intrinsic inability of CODERS to detect random responding. As the results greatly improve the later the onset of carelessness, it seems that the autoencoder needs enough attentive responses to learn the challenging task of distinguishing random responses. In other words, this is an issue of statistical power, which explains why it can be mitigated to some extent by choosing a less conservative significance level. For additional investigation, we apply traditional methods for identifying careless respondents also used by Johnson, (2005). They do not improve upon CODERS (see Appendix C.1), therefore further highlighting the difficulty of detecting random respondents in this simulation design. Coming back to CODERS, it shows similar but far less pronounced behavior for extreme responding (see the second row in Figure 8). The FNR remains decent in the early onset regime, while it is good in the baseline onset regime and excellent (close to 0) in the late onset regime.
To summarize our findings, CODERS performs well for extreme responding and near-perfectly for the two types of invariable responding. On the other hand, it struggles with random responding in this simulation design, where decent power is only achieved in case of late carelessness onset and with a not-too-conservative significance level. However, results for traditional methods (see Appendix C.1) suggest that distinguishing random responding from attentive responding may be an intrinsically difficult problem for the chosen correlation structure and marginal distributions. Overall, we find that (i) CODERS reliably discriminates between attentive and partially careless respondents, as indicated by near-zero false positive rates and low overall false negative rates, (ii) CODERS accurately estimates the location of carelessness onset, and (iii) the performance of CODERS is stable across a wide range of prevalence levels, demonstrating robustness.
Additional Simulations
We conducted additional simulation experiments to further investigate the robustness of our findings and to obtain a better understanding of possible limitations of CODERS. While details and results can be found in Appendix C, we provide a brief outline of each experiment and the most relevant findings below.
First, we vary the number of items to investigate the performance of CODERS in shorter surveys. Recall that our simulation design is based on measurements of the Big Five personality traits, with each trait being measured by 60 items (for a total of 300 items). We randomly select traits in a stepwise manner and drop all items of the selected traits, resulting in surveys of length $p\in\{240,180,120,60\}$. The overall FNR in partially careless respondents increases only slowly as survey length decreases. For $p=60$, only very few random respondents are detected, but the FNR remains decent for extreme responding and near-perfect for the invariable types of carelessness. Furthermore, the accuracy of the estimated onset of carelessness stays excellent throughout the investigated survey lengths. In summary, other than for random respondents, overall power remains adequate when using a not-too-conservative significance level (e.g., 0.5% or 1%).
[…]
Third, we consider an additional type of careless responding to investigate the performance of CODERS when the lines between attentive and careless responding are blurry. This type of carelessness, to which we refer as middling, is characterized by randomly choosing from the three middle answer categories (out of five). Given the correlation structure and marginal distributions of attentive responding (see Figure 5), there is plenty of natural variation among the central response categories, thus making middling extremely hard to distinguish. Indeed, although CODERS maintains a low FPR in fully attentive respondents, it detects only very few middling respondents (as do the traditional methods). This further highlights that CODERS requires (inconsistent) partially careless responding to be sufficiently distinct from attentive responding to successfully distinguish between them.
Fourth, we emulate a situation of temporary carelessness with two changepoints: the first constitutes a change from attentive to careless responding, and the second embodies a change back to attentive responding. We find that CODERS fails to detect such temporarily careless respondents, but the same applies to the traditional methods (except for the longstring index, which successfully captures temporary straightliners). This highlights that while CODERS is effective in detecting a single changepoint, it is simply not designed for detecting multiple changepoints.
Finally, having so far considered simulations based on personality assessments, we include a simulation inspired by a study from organizational research. This simulated survey is shorter, with $p=198$ items measuring a variety of organizationally relevant constructs, and we apply a different strategy for randomization better suited in organizational contexts. Furthermore, there exist complex relationships between (some of) the measured constructs. The FPR of CODERS is somewhat more elevated for high prevalence levels. Interestingly, the FNR is slightly lower for random responding but slightly higher for extreme responding, resulting in a comparable FNR overall. All in all, the results for CODERS remain relatively similar to those in the simulations based on personality assessments, indicating that its performance may not depend much on the institutional setting of the survey.
@@ Results
After studying the performance of CODERS in simulations inspired by the data from Johnson, (2005), we now re-examine the responses from this seminal study for partial carelessness.
@@
The data collected by Johnson, (2005) concern the five factor model of personality (Goldberg,, 1992), which assumes that variation in a measurement of personality can be explained by the so-called Big Five personality traits (factors): openness, conscientiousness, extraversion, agreeableness, and neuroticism. Each of the Big Five factors can be further decomposed into six subdomains (facets): for instance, depression and anxiety are facets of the neuroticism trait. Consequently, common Big Five instruments, such as the 240-item NEO-PI-R (Costa & McCrae,, 1992), typically measure a total of $5\times 6=30$ facets. The data set of Johnson, (2005) consists of internet-collected responses to a 300-item representation of the NEO-PI-R inventory (Goldberg,, 1999), in which each of the 30 facets is measured by 10 items on a five-point Likert scale.
@@
In general, it seems likely that survey fatigue is present in a sample of responses to an online questionnaire with a lengthy battery of 300 items. For instance, Bowling et al., (2021) estimate that if one wishes that at least 90% of participants do not respond carelessly to more than 20% of all items, one should not include more than 176 items in an online questionnaire. By applying CODERS—which is specifically designed to identify partially careless responding through a change in response behavior—in addition to the traditional screeners applied by Johnson, (2005), new and complimentary insights into the incidence and nature of (partially) careless responding may emerge.
@@
General Overview
For obtaining the autoencoder reconstruction errors in CODERS, we set the number of nodes in the bottleneck layer to 30, i.e., equal the number of facets in the Big Five model. Table 1 summarizes the results of CODERS. Depending on the significance level $\alpha$, CODERS flags between 181 (0.8%) participants for $\alpha=0.001$ and 856 (3.8%) participants for $\alpha=0.01$ as partially careless. On average, the location of the carelessness onset item for flagged participants is estimated around the 145th item, meaning that carelessness onsets on average after nearly half of the 300 items. These findings are broadly in line with the results of Bowling et al., (2021, Figure 1), whose model suggests a low probability that a representative participant starts responding carelessly at the 145th item. The earliest onset occurs at item 15, whereas the latest onset is at item 294. Figure 10 provides a more detailed visual overview of the detected onset items for the flagged participants at different significance levels.
Table 1: Results of CODERS for the data of Johnson, (2005).
Detected onset
Significance level Number of flagged respondents Minimum Average Maximum
$\alpha=0.001$ 181 (0.8%) 19 143.7 292
$\alpha=0.005$ 571 (2.5%) 19 145.8 294
$\alpha=0.01\phantom{0}$ 856 (3.8%) 15 147.9 294
[…]
Figure 10: Estimated onset of careless responding for participants flagged by CODERS in the data of Johnson, (2005).
Note. Results for different significance levels $\alpha\in\{0.01,0.05,0.1\}$ are presented in separate panels. The $x$-axis shows the indices of the participants flagged as partially carelessness by CODERS. Here, the participants are sorted in ascending order of the estimated carelessness onset item, which is in turn depicted on the $y$-axis. The dashed horizontal line represents the average location of carelessness onset across the flagged participants.
We now turn to analyzing the agreement between CODERS and the two traditional methods applied by Johnson, (2005), personal reliability and psychometric antonym. Following Johnson,, we reverse psychometric antonym scores so that low scores are associated with carelessness in both methods, and we flag respondents as careless whose personal reliability and psychometric antonym scores are lower than $0.3$ and $-0.03$, respectively. The results are shown in Table 2. Personal reliability flags 140 participants (0.6%) as careless, while psychometric antonym flags 284 respondents (1.3%).66endnote: 6We use the R package careless (Yentes & Wilhelm,, 2023) to compute personal reliability and psychometric antonym scores. The use of different software implementations (with potentially different default values) may explain differences with the numbers reported by Johnson, (2005). CODERS, personal reliability, and psychometric antonyms generally show little agreement regarding the detected careless respondents. Personal reliability and psychometric antonym scores correlate by a low value of 0.314, yielding only 11 participants that are flagged by both methods. Depending on the significance level $\alpha$, CODERS and personal reliability have between 2 and 10 flagged participants in common, while the number of jointly flagged respondents for CODERS and psychometric antonym ranges between 9 and 24. All three methods share a single flagged participant.
Table 2: Complementary findings of CODERS and traditional methods for the data of Johnson, (2005).
@@
Reliability 140 (0.6%) 2 18 10
Antonym 284 (1.3%) 9 18 24
Both 11 1 11 21
@@
It is not surprising that there is little consensus of CODERS with the two traditional methods, as they are designed for different tasks. The traditional methods aim at identifying which respondents are careless, whereas CODERS aims to detect the onset of careless responding in partially careless respondents. Hence, this disagreement indicates that (i) there may be respondents who respond carelessly throughout the survey, which CODERS cannot identify due to the absence of a changepoint (but which are found by the traditional methods), and (ii) partial carelessness may be present in the data, for which traditional methods may have low detection power (but which are found by CODERS). Moreover, CODERS uncovers novel insights on careless responding behavior, which we highlight with a selection of respondents below.
Further Insights on Individual Cases
To gain insights into the behavior of selected respondents, we analyze their per-item reconstruction errors and LSP sequences, which are shown in Figure 11. Table 3 contains additional summary statistics for these measurements before and after the estimated changepoint item (if such an item is found by CODERS). While these examples are naturally selective, similar plots for all participants flagged by CODERS or the two traditional methods can be found in our online repository at https://github.com/mwelz/carelessonset_plots.
Figure 11: Scores and changepoint item detected by CODERS in selected participants from the data of Johnson, (2005).
(a) Presumably inconsistent. Onset estimated by CODERS at item 148.
(b) Presumably inconsistent. Onset estimated by CODERS at item 246.
(c) Presumably inconsistent. Changepoint detected by CODERS at item 63.
(d) Presumably inconsistent first, then invariable. Changepoint detected by CODERS at item 58.
(e) Presumably invariable. Changepoint detected by CODERS at item 142.
(f) Presumably inconsistent throughout the survey (no changepoint item detected by CODERS).
Note. In each subplot, autoencoder reconstruction errors (RE) and LongStringPattern (LSP) sequences are shown in separate panels. The solid vertical lines indicate the identified changepoint items. Summary statistics for the corresponding participants are provided in Table 3.
Table 3: Average scores before and after the changepoint item detected by CODERS in selected participants from the data of Johnson, (2005).
@@
LSP 4.122 (1.428) 3.373 (1.464)
(a) Presumably inconsistent. Onset estimated by CODERS at item 148.
Before After
RE 0.017 (0.026) 0.070 (0.096)
LSP 2.886 (1.189) 3.236 (1.815)
(b) Presumably inconsistent. Onset estimated by CODERS at item 246.
Before After
RE 0.095 (0.119) 0.024 (0.036)
LSP 2.565 (0.917) 2.668 (0.925)
(c) Presumably inconsistent. Changepoint detected by CODERS at item 63.
@@
LSP 2.158 (0.621) 19.708 (11.027)
(d) Presumably inconsistent first, then invariable. Changepoint detected by CODERS at item 58.
@@
RE 0.018 (0.029) 0.024 (0.046)
LSP 146.000 (0.000) 23.597 (25.878)
(e) Presumably invariable. Changepoint detected by CODERS at item 142.